# The lattice packing problem in dimension 9 by Voronoi’s algorithm

### v1

@@ 7.5. Basis of minimal vectors
Up to dimension $7$ all perfect lattices have a basis of minimum vectors, which was shown theoretically by Csóka [Csó87] and confirmed by the classifications of perfect forms in those dimensions. Martinet [Mar03, Conjecture 6.6.7] conjectured that the same was true in dimension $8$, which was later confirmed by the classification [DSSV07]. He also gave a counterexample for dimension $9$, namely Coxeter’s lattice $\mathsf{A}_{9}^{2}$ [Cox51]. Our classification shows that this is in fact the only case.
@@ 7.6. Hollow perfect lattices
A lattice of dimension $n$ is hollow [MB04] if it does not have any perfect $r$-dimensional sections with the same minimum for $1<r<n$. For $n\geq 10$ perfect hollow lattices with an odd minimum always exist [MB04]. For $3\leq n\leq 8$ (the scaling of) $\mathsf{E}_{7}^{*}$ is the unique perfect hollow lattice. This left open the case for dimension $9$.
@@
All $9$-dimensional perfect lattices contain the perfect root lattice $\mathsf{A}_{2}$ as a section with the same first minimum, except Coxeter’s lattice $\mathsf{A}_{9}^{5}$. Coxeter’s lattice $\mathsf{A}_{9}^{5}$ is the unique hollow perfect lattice in dimension $9$. In particular, there are no hollow (in the sense of  [MB04]) perfect lattices with an odd minimum in dimension $9$.
@@ 7.7. The Bergé-Martinet invariant
The Bergé-Martinet invariant of a lattice $\mathcal{L}$ is defined by $\gamma^{\prime}(\mathcal{L})=\lambda_{1}(\mathcal{L})\cdot\lambda_{1}(\mathcal{L}^{*})$ [BM89]. One could interpret the Bergé-Martinet invariant as some sort of average packing density over the primal and dual lattice. This invariant has a perfectness and eutaxy theory behind it similar to Hermite’s invariant, which we will shortly recall. For more details see [Mar03, Section 3.8]. Lattices that are a local maximum of $\gamma^{\prime}$ are called dual-extreme. Lattices represented by a form $Q$ are dual extreme if and only if they are dual-perfect ($\{Qxx^{t}Q:x\in\operatorname{Min}(Q)\}\cup\{yy^{t}:y\in\operatorname{Min}(Q^{-1})\}$ span $\mathcal{S}^{n}$) and dual-eutactic (there exists a eutaxy relation $Q\left(\sum_{\pm x\in\operatorname{Min}(Q)}\lambda_{x}xx^{t}\right)Q=\sum_{\pm y\in\operatorname{Min}(Q^{-1})}\lambda_{y}yy^{t}$ where $\lambda_{x},\lambda_{y}\in\mathbb{R}_{>0}$ are strictly positive). Perfect lattices are trivially dual-perfect, and thus they are potentially a good source for dual-extreme forms.
@@
The three perfect cases attaining $\frac{16}{5}$ were already known by Martinet [Mar03] and Baril [Bar96].

### v2

@@ 7.5. Basis of minimal vectors
Up to dimension $7$ all perfect lattices have a basis of minimum vectors, which was shown theoretically by Csóka [csoka1987there] and confirmed by the classifications of perfect forms in those dimensions. Martinet [martinet, Conjecture 6.6.7] conjectured that the same was true in dimension $8$, which was later confirmed by the classification [PerfectDim8]. He also gave a counterexample for dimension $9$, namely Coxeter’s lattice $\mathsf{A}_{9}^{2}$ [coxeter1951extreme]. Our classification shows that this is in fact the only case.
@@ 7.6. Hollow perfect lattices
A lattice of dimension $n$ is hollow [martinet2004generalization] if it does not have any perfect $r$-dimensional sections with the same minimum for $1<r<n$. For $n\geq 10$ perfect hollow lattices with an odd minimum always exist [martinet2004generalization]. For $3\leq n\leq 8$ (the scaling of) $\mathsf{E}_{7}^{*}$ is the unique perfect hollow lattice. This left open the case for dimension $9$.
@@
All $9$-dimensional perfect lattices contain the perfect root lattice $\mathsf{A}_{2}$ as a section with the same first minimum, except Coxeter’s lattice $\mathsf{A}_{9}^{5}$. Coxeter’s lattice $\mathsf{A}_{9}^{5}$ is the unique hollow perfect lattice in dimension $9$. In particular, there are no hollow (in the sense of  [martinet2004generalization]) perfect lattices with an odd minimum in dimension $9$.
@@ 7.7. The Bergé-Martinet invariant
The Bergé-Martinet invariant of a lattice $\mathcal{L}$ is defined by $\gamma^{\prime}(\mathcal{L})=\lambda_{1}(\mathcal{L})\cdot\lambda_{1}(\mathcal{L}^{*})$ [berge1989probleme]. One could interpret the Bergé-Martinet invariant as some sort of average packing density over the primal and dual lattice. This invariant has a perfectness and eutaxy theory behind it similar to Hermite’s invariant, which we will shortly recall. For more details see [martinet, Section 3.8]. Lattices that are a local maximum of $\gamma^{\prime}$ are called dual-extreme. Lattices represented by a form $Q$ are dual extreme if and only if they are dual-perfect ($\{Qxx^{t}Q:x\in\operatorname{Min}(Q)\}\cup\{yy^{t}:y\in\operatorname{Min}(Q^{-1})\}$ span $\mathcal{S}^{n}$) and dual-eutactic (there exists a eutaxy relation $Q\left(\sum_{\pm x\in\operatorname{Min}(Q)}\lambda_{x}xx^{t}\right)Q=\sum_{\pm y\in\operatorname{Min}(Q^{-1})}\lambda_{y}yy^{t}$ where $\lambda_{x},\lambda_{y}\in\mathbb{R}_{>0}$ are strictly positive). Perfect lattices are trivially dual-perfect, and thus they are potentially a good source for dual-extreme forms.
@@
The three perfect cases attaining $\frac{16}{5}$ were already known by Martinet [martinet] and Baril [baril1996autour].